# Cross-domain Chinese Sentence Pattern Parsing

### 27 Feb 2024 v2

@@ 2.1 Cross-domain Syntactic Parsing
Syntactic parsing (Collins, 1997; McClosky et al., 2006; Finkel et al., 2008; Zhu et al., 2013; Dyer et al., 2015; Dozat and Manning, 2016; Stern et al., 2017; Gaddy et al., 2018; Guo et al., 2023) is the automatic analysis of the syntactic structure of natural language. It is a long-standing traditional NLP task, has evolved over the years. There has been work considering cross-domain constituent parsing with parser combinations (McClosky et al., 2010). Furthermore, cross-domain parsing has also been investigated on other grammar formalisms, specifically dependency syntax (Blodgett et al., 2018; Li et al., 2019; Rotman and Reichart, 2019). However, unsupervised or weakly supervised open domain parsers remain a challenging task (Yang et al., 2022). There is limited research on cross-domain constituency parsing (McClosky et al., 2010; Fried et al., 2019; Li et al., 2023a), and even less on cross-domain SPS parsing (Xie et al., 2022).
@@ 2.3 Self-training
Self training (Yarowsky, 1995; McClosky et al., 2006; Yu et al., 2015; Ramponi and Plank, 2020; Guo et al., 2023) has emerged as a promising approach for domain adaptation, applied to tasks such as classification (Dong and Schäfer, 2011; Ye et al., 2020), sequence labeling (Wang et al., 2021), dependency parsing (Yu et al., 2015; Rotman and Reichart, 2019; Guo et al., 2023), sequence generation (He et al., 2019), and QA (Sachan and Xing, 2018) and so on. Recently, some researchers have also applied it to cross-domain Cross-domain Constituency Parsing (Li et al., 2023a). In this work, we focus on unsupervised cross-domain SPS parsing, investigating the LLM-enhanced self-training guided by rules.
@@ 2.4 LLMs Parsing
LLMs have been successfully applied to various NLP tasks (Brown et al., 2020; He et al., 2023; Mysore et al., 2023; Zhang et al., 2023; Pangakis et al., 2023; Ashok and Lipton, 2023; Bai et al., 2023), including commonsense reasoning (Zhang et al., 2022), text summarization (Goyal et al., 2022; Chen et al., 2023), and massive multitask language understanding (Hendrycks et al., 2021). Although LLMs have shown limitations in structure extraction tasks (Qin et al., 2023), such as cross-domain constituency parsing, their generative capabilities can be utilized to provide a raw corpus for self-training SPS parsing (Li et al., 2023a).
Due to the hallucination issues of LLMs, we implemented syntactic rules to standardize the structure of generated sentences. Previous research, as demonstrated by Yang et al. (2022), has shown a close relationship between grammatical rules and the effectiveness of constituency parsing. Meanwhile, Wang et al. (2023) successfully utilized grammatical rules for generating cross-domain data. Our work integrates LLMs into the iterative process of self-training (Li et al., 2023a), utilizing syntactic rules to generate raw corpora.
@@ 3.1 Berkeley Neural Parser
We utilize the Berkeley Neural Parser (Kitaev and Klein, 2018) as the base for our approach. It is chart-based and incorporates a self-attentive encoder along with a chart-based decoder. It uniquely identifies span labels using only the local span representations without directly modeling the structure of the output tree.
Due to the integration of pre-trained language models, the Berkeley Neural Parser inherently possesses cross-domain constituency parsing capabilities. This enables the parser to be trained on the source domain and then directly applied to parse trees for the target domain. We use this direct model transfer as a baseline for comparison.
@@ 3.2 LLM-enhanced Self-training
The primary objective of self-training is to generate high-quality training instances for the target domain, which are then utilized to train the target domain model. Vanilla self-training for cross-domain parsing refers to an iterative process aimed at training a parser for a target domain.
We dynamically incorporate the LLM-enhanced method into the self-training process to improve the quality and quantity of raw corpora during iterations. In each iteration, we extract syntactic rules from the training data and input them into the LLMs with a few example sentences, to generate raw corpora for the target domain. As the iterative process progresses, the parser’s capability gradually improves. As shown in the Figure 2, the process specifically includes the following steps:
@@ 3.3 Rule-based Instances Selection Criteria
To filter higher-quality pseudo-data from the raw corpus generated by the LLM, we propose rule-based instance selection criteria. Unlike previous self-training selection criteria that focus solely on the task, this criterion considers the syntactic features of the corpus generated by the LLMs to ensure that the filtered pseudo-data are suitable for cross-domain parsing self-training. The feasibility of grammar-rule-based selection criteria is also supported by Yang et al. (2022) and Wang et al. (2023).
We employ Equation (1) to assess the distance between the training set and individual instances:
@@
where $S$ represents the source domain dataset in the case of the foundational selection criteria. After the incorporation of a rule-based method, $S$ becomes the target domain set transformed by rules. The candidate instances are denoted as $c\in C$ (the candidate set).
We then select the top $K$ candidate sentences that are closest to the source domain set as extended training instances during the self-training process. This approach ensures the selection of the most relevant instances, thereby achieving a gradual adaptation to the target domain.
Our instance selection process encompasses three levels of criteria: token, confidence, and syntactic rule. Additionally, we have attempted to integrate criteria based on confidence and syntactic rules to establish a more effective standard, aimed at better filtering instances that adapt to the target domain.
@@
3.4 LLMs Prompt
To generate higher-quality sentences that are closer to the target domain, we introduce an LLMs prompt. This prompt includes an introduction to the sentence pattern tags, syntactic rules, and examples from the target domain. During the generation, we prepare the following parameters: 1) N syntactic rules extracted from the treebank, 2) M sentences sampled from the target domain, and 3) length constraints L1 $\sim$ L2 for the generated sentence to ensure that the length of the generated sentences matches the length distribution of target domain sentences.
Syntactic Rules. We determine the value of N by sampling from the distribution of sentence lengths in the treebank and extracting syntactic rules from it. Specifically, we introduce a Gaussian distribution for the average length of source domain data, denoted as N $=N(avg\_len,6)$, to obtain N syntactic rules.
Sampled Sentences. Due to resource constraints and minimal performance differences, based on preliminary experiments, we opt to randomly sample 5 sentences from existing target domain data as input.
Length Constraints. To ensure that the length distribution of the generated sentences is more closely aligned with that of the gold standard target domain sentences, we employ another normal distribution, N $=N(avg\_len,3)$ to sample two values, L1 and L2. These values define the limits for the length of the sentences to be generated.
The example of the LLM prompt is illustrated in Figure 3.

### 08 Apr 2024 v3

@@ 2.1 Cross-domain Syntactic Parsing
Syntactic parsing (Collins, 1997; McClosky et al., 2006; Finkel et al., 2008; Zhu et al., 2013; Dyer et al., 2015; Dozat and Manning, 2016; Stern et al., 2017; Gaddy et al., 2018; Guo et al., 2023) is the automatic analysis of the syntactic structure of natural language. It is a long-standing traditional NLP task, has evolved over the years. Research has explored the application of parser combinations (McClosky et al., 2010) and LLM-enhanced (Li et al., 2023a) for cross-domain constituent parsing. Furthermore, cross-domain parsing has also been investigated on other grammar formalisms, specifically dependency syntax (Blodgett et al., 2018; Li et al., 2019; Rotman and Reichart, 2019). For SPS parsing, the work by Xie et al. (2022) engaged with the issue of cross-domain applicability but fell short of providing a comprehensive solution. Their methodology, which is based on rule-based mapping, specifically designs mapping rules from constituent trees to SPS trees for Penn Chinese Treebank (CTB). However, this approach is not adaptable for application across various domains, limiting its generalizability.
@@ 2.3 Self-training
Self training (Yarowsky, 1995; McClosky et al., 2006; Yu et al., 2015; Ramponi and Plank, 2020; Guo et al., 2023) is a semi-supervised learning technique mainly utilized in scenarios with a limited amount of labeled data complemented by a substantial volume of unlabeled data. This method initiates with the training of a foundational model using the available labeled data, and then iteratively generates pseudo-data with high confidence, incrementally enlarging the dataset for model training. This process aims to enhance the model’s performance and generalization capability in the context of scarce labeled data. Self-training has been extensively applied across various NLP tasks (Dong and Schäfer, 2011; Ye et al., 2020; Wang et al., 2021; Yu et al., 2015; Rotman and Reichart, 2019; Guo et al., 2023; He et al., 2019; Sachan and Xing, 2018; Li et al., 2023a). In this work, we focus on unsupervised cross-domain SPS parsing, investigating the LLM-enhanced self-training guided by rules.
@@ 2.4 LLMs Parsing
Due to the remarkable generative capabilities of LLMs, they have achieved success in numerous NLP tasks (Brown et al., 2020; He et al., 2023; Mysore et al., 2023; Zhang et al., 2023; Pangakis et al., 2023; Ashok and Lipton, 2023; Bai et al., 2023), including commonsense reasoning (Zhang et al., 2022), text summarization (Goyal et al., 2022; Chen et al., 2023), and massive multitask language understanding (Hendrycks et al., 2021). Research has combined LLMs with constituency parsing, demonstrating the efficacy of LLMs in syntactic parsing. Li et al. (2023a) employed LLMs for data augmentation, generating raw corpora across five domains and dynamically incorporating them into the training process. Bai et al. (2023) assessed the performance of LLMs under zero-shot, few-shot, and full-training settings, evaluating both in-domain and out-of-domain performance.
@@ 3 Method
3.1 Rule-based Mapping
Research on constituency parsing is relatively mature, with parsers achieving high accuracy. Consequently, Xie et al. (2022) designed rules for the CTB to convert constituent trees into SPS trees. We first generate constituent trees for our testing set with CoreNLP, then apply rule-based mapping to obtain SPS trees for comparative analysis.
@@ 3.2 Berkeley Neural Parser
The Berkeley Neural Parser (Kitaev and Klein, 2018) is a state-of-the-art tool in syntactic analysis, leveraging neural networks to accurately generate parse trees from sentences. It is distinguished by its ability to adapt across languages and domains, thanks to its integration with pre-trained language models. Its architecture, which combines a self-attentive encoder with a chart-based decoder, allows for efficient and reliable parsing. This has made the parser a preferred choice for syntactic parsing tasks among researchers and practitioners alike. Based on the bert-base-chinese model, we continue training with textbook data and news data after rule-based mapping, serving as our foundational baseline.
@@ 3.3 LLM-enhanced Self-training
To improve data diversity, we supply LLMs with syntactic rules and example sentences extracted from the target domain, facilitating the generation of candidate raw corpora for each iteration of self-training. As the iterative process progresses, the parser’s capability gradually improves. As shown in the Figure 2, the process specifically includes the following steps:
@@
To generate higher-quality sentences that more closely resemble the target domain, we constrained the prompts for LLMs from three aspects: syntactic rules, example sentences, and length. We introduced a Gaussian distribution for the average length of source domain data and randomly sampled values at each request to constrain the length of the generated sentences, as well as the number of required syntactic rules.
The example of the LLM prompt is illustrated in Figure 3.
To verify whether LLMs genuinely adhere to our constraints during data generation, we randomly extracted 20 sentences of data from the raw corpus generated by LLMs. Then manual annotation and matched them with the syntactic rules provided in the prompt. We found that the syntactic structure of the generated sentences matched the provided syntactic rules 69.3% of the time. Furthermore, when the prompt includes uncommon labels such as $ind$ (independent word), LLMs tend to generate sentences with the $ind$ label.
@@ 3.4 Rule-based Instances Selection Criteria
Yang et al. (2022) and Wang et al. (2023) have confirmed the feasibility of grammar-rule-based selection criteria. To filter higher-quality pseudo-data from the raw corpus generated by the LLM, we propose mapping-rule-based instance selection criteria. Unlike previous self-training selection criteria that focus on grammar rules extracted from the source domain (Li et al., 2023a), this criterion considers the selection based on syntactic rules from the target domain.
We first get target domain SPS trees converted via mapping rules. Differently from previously calculating the distance between pseudo-data and training data, we calculate the distance between pseudo-data and the converted target domain data to filter pseudo-data suitable for cross-domain parsing self-training.
The Jensen-Shannon (JS) divergence is a symmetric and finite measure of similarity between two probability distributions, often used to quantify the difference in information content or to assess the similarity between datasets. We employ the JS divergence shown in Equation (1) to assess the similarity between the target data converted by mapping rules and individual instances:
@@
where $S$ represents the target domain dataset transformed by rules, and $C$ is the candidate set, as well as the candidate instances are denoted as $c\in C$.
We then select the top $K$ candidate sentences as shown in Equation (2). These sentences possess syntactic structures more closely aligned with the converted target domain data and serve as an extension of the training set in subsequent iterations of self-training to achieve a gradual adaptation to the target domain.
Our criteria has three levels: token, confidence, and syntactic rule. Additionally, we considered the integration of criteria based on confidence and syntactic rules to establish a more effective standard, aimed at better filtering instances that adapt to the target domain, whick is also supported by Li et al. (2023a).